A quadrilateral enclosure has sides measuring 89.7 ft, 93.75ft, 91.279 ft and 88.52 ft. Use the correct number of significant digits to find the perimeter of the enclosure.

The spinner below is spun twice. Find the probability of the outcome. Find the P(a number less than 3 or a number greater thaN 5. sPINNER NUMBERED 1,2,3,4,5,6,7,8,. tHANK YOU SOO MUCH FOR YOUR TIME. CATCH YOU ON FRIDAY WITH MY CARD.

Answer by jim_thompson5910(35256)

Perimeter: Sum of all 4 sides

P = 49.7+93.75+91.279+88.52

P = 323.249

-------------------------------------------------------

There are 2 numbers less than 3 (1 and 2) and 3 numbers greater than 5 (6,7,8) out of 8 total. So the probability is


P(A number less than 3 or greater than 5) = (2+3)/8 = 5/8